The question gives that the diameter of the cylinder is 54 inches and the height is 159 inches
find the surface area of the cylinder

Answer by lwsshak3(11628)

The question gives that the diameter of the cylinder is 54 inches and the height is 159 inches
find the surface area of the cylinder.
***
surface area of cylinder=area of 2 circular ends+lateral surface area
area of 2 circular ends=2*π*54^2/4=1458π sq in
lateral surface area=circumference*height=π*54*159=8586π sq in
total surface area of cylinder=1458π+8586π=10044π sq in